solve 8X^2=800. Is the answer X=5(radical)2/2 or X=5+or-(radical)2/2???
Answer by stanbon(75887)

solve 8X^2=800
Divide both sides by 8 to get:
x^2=100
x^2-100 = 0
Factor:
(x-10)(x+10) = 0
x = 10 or x = -10
======================
cheers,
Stan H.